### v1

@@ Abstract
Military personnel and security agents often face significant physical risks during conflict and engagement situations, particularly in urban operations. Ensuring the rapid and accurate communication of incidents involving injuries is crucial for the timely execution of rescue operations. This article presents research conducted under the scope of the Brazilian Navy’s “Soldier of the Future” project, focusing on the development of a Casualty Detection System to identify injuries that could incapacitate a soldier and lead to severe blood loss. The study specifically addresses the detection of soldier falls, which may indicate critical injuries such as hypovolemic hemorrhagic shock. To generate the publicly available dataset, we used smartwatches and smartphones as wearable devices to collect inertial data from soldiers during various activities, including simulated falls. The data were used to train 1D Convolutional Neural Networks (CNN1D) with the objective of accurately classifying falls that could result from life-threatening injuries. We explored different sensor placements—on the wrists and near the center of mass—and various approaches to using inertial variables, including linear and angular accelerations. The neural network models were optimized using Bayesian techniques to enhance their performance. The best-performing model and its results, discussed in this article, contribute to the advancement of automated systems for monitoring soldier safety and improving response times in engagement scenarios.
Index Terms:
Fall Detection, Machine Learning, Artificial Neural Networks.
[…]
=-21pt
I Introduction
@@
The same procedure was applied to the angular acceleration of the gyroscope, and we experimented with an approach that combined the linear and angular acceleration data. The aim was to determine which of these approaches in using inertial variables would be more efficient. These different approaches were organized into multiples experimental pipelines, each identified by an acronym (see Section III-C3).
@@
This paper is divided into five sections. Section I corresponds to the already presented introduction, while Section II addresses related work, covering relevant research that underpinned this study. Section III is dedicated to materials and methods, describing the IPqM-FALL dataset, data collection methods, data preparation and analysis, the architectures of the neural networks used, and the Bayesian optimization process. Section IV presents the results obtained, accompanied by their discussion. Finally, Section V on the conclusion provides the final considerations on the research conducted.
II Related Work
@@
III Materials and Methods
This section is dedicated to detailing the methods employed in data collection, the creation of the IPqM-Fall dataset, as well as data preparation and the selection of neural network architectures. This section is organized as follows. Section III-A explores the operation of the casualty detection system, explaining in detail how the data is monitored and analyzed to identify casualty occurrences. Section III-B describes the data collection procedures and the structure of the IPqM-Fall dataset. Section III-C discusses the selection and configuration of neural network architectures, also addressing Bayesian hyperparameter optimization and the different testing pipelines used to validate the models.
III-A System Design
@@
III-B IPqM Fall Dataset
In this section, we provide an overview of the IPqM-FALL dataset, detailing the methodology for data collection and organization for fall detection. The aim was to create a comprehensive database for analyzing everyday activities, military operations, and falls, using inertial variables collected by wearable devices and smartphones. The following sections address the devices used, and the variables collected (Section III-B1), the data collection methodology (Section III-B2), the voluntary participants (Section III-B3), the collected Activities (Section III-B4), and the organization of the dataset (Section III-B5). Each section offers insight into the methods employed to ensure the integrity and utility of the data for training and validating machine learning models for fall detection.
III-B1 Devices Used and Collected Variables
@@
III-B2 Data Collection
The collection method was designed to capture data from multiple individuals simultaneously. Each volunteer military personnel used three devices during the activities: an Android smartphone in the chest pocket of the military uniform, along with two WearOS smartwatches, one on the left wrist and another on the right wrist (Figure 3). Custom sampling apps were developed and installed on both the smartphones and smartwatches, along with an instructor app installed on an Android tablet used by the experiment coordinator. Concretely, two sampling apps were developed: one for mobile devices and another for wearable platforms.
Figure 3: Position in the body of the devices used.
@@ III-B3 Voluntary Participants
Military personnel of both sexes and various ranks were selected as shown in Table I. The selection of personnel from different ranks aims to ensure a diversity of experiences and skills, providing a more representative sample of the military population.
TABLE I: Age, rank, sex and anthropometric information of volunteer military personnel.
@@
III-B4 Collected Activities
As mentioned earlier in Section III-B2, three distinct types of activities were collected: Activities of Daily Living (ADL), which are basic tasks performed in everyday life; Military Operations Activities, which involve tasks performed in military operations and service involving firearms; and Fall Activities, which include situations involving loss of balance and contact of the body with the ground. These activities were selected to cover a wide range of situations experienced by a soldier, including those that might be confused with falling activities, with the aim of obtaining a diverse dataset for training neural networks or other machine learning techniques.
The ADL data collected were as follows: 1) standing, 2) walking, 3) running, 4) jumping, 5) sitting in a chair, 6) standing up from a chair, 7) walking uphill, 8) walking downhill, 9) running uphill, 10) running downhill, 11) walking upstairs, 12) walking downstairs, 13) jumping upstairs one step. The number of repetitions and the duration of these activities, along with instructions, are detailed in Table II.
TABLE II: Activities of Daily Living.
@@ III-B4 Collected Activities
ADL_1 Standing 1 3 min Standing upright, performing subtle movements
ADL_2 Walking 1 3 min Walking throughout the activity duration
ADL_3 Running 3 30 sec Running throughout the activity duration
ADL_4 Jumping 3 30 sec Standing, performing short jumps every 2 seconds
ADL_5 Stair climbing 3 10 sec Walking upstairs
ADL_6 Stair descending 3 10 sec Walking downstairs
ADL_7 Sitting in chair 6 5 sec Starts standing, transitions to sitting in a chair
ADL_8 Standing from chair 6 5 sec Starts sitting in a chair, transitions to standing
ADL_11 Uphill walking 3 30 sec Walking uphill
ADL_12 Downhill walking 3 30 sec Walking downhill
ADL_13 Uphill running 3 15 sec Running uphill
ADL_14 Downhill running 3 15 sec Running downhill
ADL_15 Stair hopping 3 10 sec Climbing the stairs and jumping the space one step at a time
Activities ADL_1, ADL_4, ADL_5, ADL_6, ADL_11, ADL_12, ADL_13, and ADL_14 were also performed while carrying a rifle, considered as military operation activities with the same number and duration of repetitions indicated in Table II. Activities ADL_13 and ADL_14 with a rifle are not conventional running, but rather speed engagement activities.
The Military Operation Activities (MO) were performed while carrying a rifle, except for the crawling activity. The collected activities were as follows: 1) Sweep with walking, 2) Sweep with quick engagement, 3) Transition from standing to kneeling shooting position, 4) Transition from walking to kneeling shooting position, 5) Transition from running to kneeling shooting position, 6) Transition from standing to prone shooting position, 7) Transition from walking to prone shooting position, 8) Transition from running to prone shooting position, 9) Crawl. The number and duration of the activities, as well as the descriptions, are detailed in Table III.
TABLE III: Activities of Military Operations.
@@
OM_1 Sweep (Walking) 1 3 min Walking while sweeping (pointing the rifle at a coverage angle of approximately 120 degrees)
OM_2 Sweep (Quick Engagement) 3 45 sec Quick Engaging while sweeping (pointing the rifle at a coverage angle of approximately 120 degrees)
OM_3 Kneeling Shooting Position (Standing) 3 10 sec Starts standing, transitions to kneeling shooting position
OM_4 Kneeling Shooting Position (Walking) 3 1 min Starts walking, transitions to kneeling shooting position
OM_5 Kneeling Shooting Position (Running) 3 45 sec Starts running, transitions to kneeling shooting position
OM_6 Prone Shooting Position (Standing) 3 10 sec Starts standing, transitions to prone shooting position
OM_7 Prone Shooting Position (Walking) 3 1 min Starts walking, transitions to prone shooting position
OM_8 Prone Shooting Position (Running) 3 45 sec Starts running, transitions to prone shooting position
OM_9 Crawl 3 50 sec Start in the crawling position and perform the Crawl activity for the duration of the activity.
The Falling Activities were conducted in two ways, with rifle and without rifle, as follows: 1) forward fall ending lying on the back, 2) forward fall ending lying face down, 3) backward fall, 4) right side fall, 5) left side fall. The side falls were alternated, with one group of volunteer military personnel performing the left side fall while another group performed the right side fall. The quantity, duration, and descriptions of the activities are detailed in Table IV.
TABLE IV: Fall Activities. The rifle falls were performed with the same number of repetitions and duration.
@@
FALL_1 Frontal Fall (Supine) 6 10 sec Falling forward from standing position to lying supine
FALL_2 Frontal Fall (Prone) 6 10 sec Falling forward from standing position to lying prone
FALL_3 Backward Fall 6 10 sec Falling backward from standing position to lying supine
FALL_5 Lateral Fall - Right 6 10 sec Falling to the right from standing position
FALL_6 Lateral Fall - Left 6 10 sec Falling to the left from standing position
@@
III-B5 Dataset Organization
The dataset is organized as follows: there are 15 directories corresponding to each volunteer military personnel. Within each directory, there is a file named README and three subdirectories named chest, left, and right, which contain data collected by the smartphone (on the chest) and by the smartwatches on the left and right wrists. The README file contains information about the volunteer military personnel, which was recorded in Table I, including the number of activities and repetitions performed by the volunteer. Each subdirectory contains the two files in CSV format, as described below:
@@ III-B5 Dataset Organization
III-C Construction of the Fall Detection Neural Network Model
This section discusses the steps involved in constructing the model, including data preprocessing, network architecture selection, and the training process. Additionally, it addresses key considerations such as feature extraction, model optimization, and the process of labeling.
III-C1 Choice of Neural Network Architectures
@@ III-C2 Feature Extraction
III-C3 Data Analysis and Preparation
@@
Figure 6: Fall activities recorded by the left smartwatch of ID 3: Fall_1, Fall_2, Fall_3, and Fall_5, respectively.
Figure 7: Fall activities recorded by the left smartwatch of ID 3: ADL_1, ADL_2, ADL_3, ADL_4, ADL_5 and ADL_7, respectively.
There is a difference in the magnitude of activities that involve walking and running with and without a rifle (ADL_2, ADL_3, OM_1, and OM_2), with the magnitude of activities with a rifle being lower (Figure 8). This is because walking with a rifle is an engagement walk, while running is a high-speed engagement activity, not exactly a walk or a run. Among the activities that involve falls with and without a rifle, the patterns are similar, but in many cases, the magnitude of falls with a rifle is lower (Figure 8). Generally, the fall with or without the use of a rifle usually occurs in the first five seconds, followed by a stable pattern with an average acceleration magnitude of approximately 10 in the final five seconds, corresponding to the inactivity of the fallen body on the ground. Thus, the time window defined for fall detection is five seconds, corresponding, on average, to 450 observation points for the smartwatch and 1025 observation points for the smartphone. For both linear and angular acceleration, these averages are similar due to the data collection frequencies being close.
@@
Figure 8: Fall activities recorded by the left smartwatch of ID 3: ADL_12, ADL_12_withRifle, Fall_3 and Fall_3_withRifle, respectively.
Activities with data collected during trials of ten seconds or more, except for transition activities and ADL_5, ADL_6, and ADL_15 activities related to going up and down stairs, exhibit a similar pattern in the magnitude variation over the activity duration. This allowed these activities to be segmented into more than one five-second set. For activities involving stairs, the set size was defined using only the first five seconds of the activity, discarding the rest of the collected data, as the staircase where the activities were collected was short, with only ten steps, and in most of the collected activities, after five seconds, the volunteer military personnel had already completed the ascent or descent of the stairs.
@@
In the transition activities to the shooting position (OM_3 to OM_8), such as the OM_4 activity, which is the transition from walking to the kneeling shooting position, the point of interest is the pattern caused by the state change, to verify with the neural network if, in any way, the activity is confused with a fall activity. Therefore, in these activities, only one five-second observation vector (480 and 1025) was selected and represents the moment of state change from one activity to another. In this way, the observation point where the highest magnitude peak occurs was identified, and the 240 observations before the peak and the 240 observations after the peak were selected to construct the data vector presented to the neural network, except for cases where it was not possible to distribute equally.
Fall activities and transition activities to the prone shooting position (OM_6, OM_7, and OM_8) exhibit similar patterns (Figure 9 because both involve the body hitting the ground, which can potentially confuse the neural network. However, as the first stage of the Soldier of the Future project aims to detect falls for subsequent classification if the fall occurred due to injuries or not, an adopted approach was to consider these transition activities as falls. Each experimental pipeline was executed with labels considering these activities as falls and, in another approach, considering them as non-fall activities.
Figure 9: Activities recorded by the left smartwatch of ID 3: Fall_3 and OM_6, respectively.
@@
For the execution of all experimental pipelines (Section III-C4), the pre-processed data for training the neural networks was divided into three sets: training, testing, and validation. The test and validation sets each correspond to 20% of the samples, while the training set corresponds to 60% of the samples. Bayesian optimization of the hyperparameters (Section III-C5) was performed using the training and validation data. The test set was not presented to the network during optimization or subsequent training, and was used only for making predictions and evaluating the models’ generalization performance.
@@
III-C4 Experimental Pipelines and Labeling Approach

### v2

@@ Abstract
Military personnel and security agents often face significant physical risks during conflict and engagement situations, particularly in urban operations. Ensuring the rapid and accurate communication of incidents involving injuries is crucial for the timely execution of rescue operations. This article presents research conducted under the scope of the Brazilian Navy’s “Soldier of the Future” project, focusing on the development of a Casualty Detection System to identify injuries that could incapacitate a soldier and lead to severe blood loss. The study specifically addresses the detection of soldier falls, which may indicate critical injuries such as hypovolemic hemorrhagic shock. To generate the publicly available dataset, we used smartwatches and smartphones as wearable devices to collect inertial data from soldiers during various activities, including simulated falls. The data were used to train 1D Convolutional Neural Networks (CNN1D) with the objective of accurately classifying falls that could result from life-threatening injuries. We explored different sensor placements (on the wrists and near the center of mass) and various approaches to using inertial variables, including linear and angular accelerations. The neural network models were optimized using Bayesian techniques to enhance their performance. The best-performing model and its results, discussed in this article, contribute to the advancement of automated systems for monitoring soldier safety and improving response times in engagement scenarios.
1 Introduction
@@
The same procedure was applied to the angular acceleration of the gyroscope, and we experimented with an approach that combined the linear and angular acceleration data. The aim was to determine which of these approaches in using inertial variables would be more efficient. These different approaches were organized into multiples experimental pipelines, each identified by an acronym (see Section 3.3.3).
@@
This paper is divided into five sections. Section 1 corresponds to the already presented introduction, while Section 2 addresses related work, covering relevant research that underpinned this study. Section 3 is dedicated to materials and methods, describing the IPqM-FALL dataset, data collection methods, data preparation and analysis, the architectures of the neural networks used, and the Bayesian optimization process. Section 4 presents the results obtained, accompanied by their discussion. Finally, Section 5 on the conclusion provides the final considerations on the research conducted.
2 Related Work
@@
3 Materials and Methods
This section is dedicated to detailing the methods employed in data collection, the creation of the IPqM-Fall dataset, as well as data preparation and the selection of neural network architectures. This section is organized as follows. Section 3.1 explores the operation of the casualty detection system, explaining in detail how the data is monitored and analyzed to identify casualty occurrences. Section 3.2 describes the data collection procedures and the structure of the IPqM-Fall dataset. Section 3.3 discusses the selection and configuration of neural network architectures, also addressing Bayesian hyperparameter optimization and the different testing pipelines used to validate the models.
3.1 System Design
@@
3.2 IPqM Fall Dataset
In this section, we provide an overview of the IPqM-FALL dataset, detailing the methodology for data collection and organization for fall detection. The aim was to create a comprehensive database for analyzing everyday activities, military operations, and falls, using inertial variables collected by wearable devices and smartphones. The following sections address the devices used, and the variables collected (Section 3.2.1), the data collection methodology (Section 3.2.2), the voluntary participants (Section 3.2.3), the collected Activities (Section 3.2.4), and the organization of the dataset (Section 3.2.5). Each section offers insight into the methods employed to ensure the integrity and utility of the data for training and validating machine learning models for fall detection.
3.2.1 Devices Used and Collected Variables
@@
3.2.2 Data Collection
The data collection method was designed to capture data from multiple individuals simultaneously. Each volunteer military personnel used three devices during the activities: an Android smartphone in the chest pocket of the military uniform, along with two WearOS smartwatches, one on the left wrist and another on the right wrist (Figure 3). Custom sampling apps were developed and installed on both the smartphones and smartwatches, along with an instructor app installed on an Android tablet used by the experiment coordinator. Concretely, two sampling apps were developed: one for mobile devices and another for wearable platforms.
Figure 3: Positions of the devices used in the simulated activities.
@@ 3.2.3 Voluntary Participants
Military personnel of both sexes and various ranks were selected as shown in Table 1. The selection of personnel from different ranks aims to ensure a diversity of experiences and skills, providing a more representative sample of the military population.
Table 1: Age, rank, sex and anthropometric information of volunteer military personnel.
@@
3.2.4 Collected Activities
As mentioned earlier in Section 3.2.2, three distinct types of activities were collected: Activities of Daily Living (ADL), which are basic tasks performed in everyday life; Military Operations Activities, which involve tasks performed in military operations and service involving firearms; and Fall Activities, which include situations involving loss of balance and contact of the body with the ground. These activities were selected to cover a wide range of situations experienced by a soldier, including those that might be confused with falling activities, with the aim of obtaining a diverse dataset for training neural networks or other machine learning techniques.
The ADL data collected were as follows: 1) standing, 2) walking, 3) running, 4) jumping, 5) sitting in a chair, 6) standing up from a chair, 7) walking uphill, 8) walking downhill, 9) running uphill, 10) running downhill, 11) walking upstairs, 12) walking downstairs, 13) jumping upstairs one step. The number of repetitions and the duration of these activities, along with instructions, are detailed in Table 2111The code numbering lacks $\text{ADL}_{9}$ and $\text{ADL}_{10}$ in Table 2 because these simulations were not executed later, due to operational constraints. The same goes for $\text{FALL}_{4}$ in Table 4..
Table 2: Activities of Daily Living.
@@ 3.2.4 Collected Activities
$\text{ADL}_{1}$ Standing 1 3 min Standing upright, performing subtle movements
$\text{ADL}_{2}$ Walking 1 3 min Walking throughout the activity duration
$\text{ADL}_{3}$ Running 3 30 sec Running throughout the activity duration
$\text{ADL}_{4}$ Jumping 3 30 sec Standing, performing short jumps every 2 seconds
$\text{ADL}_{5}$ Stair climbing 3 10 sec Walking upstairs
$\text{ADL}_{6}$ Stair descending 3 10 sec Walking downstairs
$\text{ADL}_{7}$ Sitting in chair 6 5 sec Starts standing, transitions to sitting in a chair
$\text{ADL}_{8}$ Standing from chair 6 5 sec Starts sitting in a chair, transitions to standing
$\text{ADL}_{11}$ Uphill walking 3 30 sec Walking uphill
$\text{ADL}_{12}$ Downhill walking 3 30 sec Walking downhill
$\text{ADL}_{13}$ Uphill running 3 15 sec Running uphill
$\text{ADL}_{14}$ Downhill running 3 15 sec Running downhill
$\text{ADL}_{15}$ Stair hopping 3 10 sec Climbing the stairs and jumping the space one step at a time
Activities $\text{ADL}_{1}$, $\text{ADL}_{4}$, $\text{ADL}_{5}$, $\text{ADL}_{6}$, $\text{ADL}_{11}$, $\text{ADL}_{12}$, $\text{ADL}_{13}$, and $\text{ADL}_{14}$ were also performed while carrying a rifle, considered as military operation activities with the same number and duration of repetitions indicated in Table 2. Activities $\text{ADL}_{13}$ and $\text{ADL}_{14}$ with a rifle are not conventional running, but rather speed engagement activities.
The Military Operation Activities (MO) were performed while carrying a rifle, except for the crawling activity. The collected activities were as follows: 1) Sweep with walking, 2) Sweep with quick engagement, 3) Transition from standing to kneeling shooting position, 4) Transition from walking to kneeling shooting position, 5) Transition from running to kneeling shooting position, 6) Transition from standing to prone shooting position, 7) Transition from walking to prone shooting position, 8) Transition from running to prone shooting position, 9) Crawl. The number and duration of the activities, as well as the descriptions, are detailed in Table 3.
Table 3: Activities of Military Operations.
@@
$\text{OM}_{1}$ Sweep (Walking) 1 3 min Walking while sweeping (pointing the rifle at a coverage angle of approximately 120 degrees)
$\text{OM}_{2}$ Sweep (Quick Engagement) 3 45 sec Quick Engaging while sweeping (pointing the rifle at a coverage angle of approximately 120 degrees)
$\text{OM}_{3}$ Kneeling Shooting Position (Standing) 3 10 sec Starts standing, transitions to kneeling shooting position
$\text{OM}_{4}$ Kneeling Shooting Position (Walking) 3 1 min Starts walking, transitions to kneeling shooting position
$\text{OM}_{5}$ Kneeling Shooting Position (Running) 3 45 sec Starts running, transitions to kneeling shooting position
$\text{OM}_{6}$ Prone Shooting Position (Standing) 3 10 sec Starts standing, transitions to prone shooting position
$\text{OM}_{7}$ Prone Shooting Position (Walking) 3 1 min Starts walking, transitions to prone shooting position
$\text{OM}_{8}$ Prone Shooting Position (Running) 3 45 sec Starts running, transitions to prone shooting position
$\text{OM}_{9}$ Crawl 3 50 sec Start in the crawling position and perform the Crawl activity for the duration of the activity.
The Falling Activities were conducted in two ways, with rifle and without rifle, as follows: 1) forward fall ending lying on the back, 2) forward fall ending lying face down, 3) backward fall, 4) right side fall, 5) left side fall. The side falls were alternated, with one group of volunteer military personnel performing the left side fall while another group performed the right side fall. The quantity, duration, and descriptions of the activities are detailed in Table 4.
Table 4: Fall Activities. The rifle falls were performed with the same number of repetitions and duration.
@@
$\text{FALL}_{1}$ Frontal Fall (Supine) 6 10 sec Falling forward from standing position to lying supine
$\text{FALL}_{2}$ Frontal Fall (Prone) 6 10 sec Falling forward from standing position to lying prone
$\text{FALL}_{3}$ Backward Fall 6 10 sec Falling backward from standing position to lying supine
$\text{FALL}_{5}$ Lateral Fall - Right 6 10 sec Falling to the right from standing position
$\text{FALL}_{6}$ Lateral Fall - Left 6 10 sec Falling to the left from standing position
@@
3.2.5 Dataset Organization
The dataset is organized as follows: there are 15 directories corresponding to each volunteer military personnel. Within each directory, there is a file named README and three subdirectories named chest, left, and right, which contain data collected by the smartphone (on the chest) and by the smartwatches on the left and right wrists. The README file contains information about the volunteer military personnel, which was recorded in Table 1, including the number of activities and repetitions performed by the volunteer. Each subdirectory contains the two files in CSV format, as described below:
@@ 3.2.5 Dataset Organization
3.3 Construction of the Fall Detection Neural Network Models
This section discusses the steps involved in constructing the fall detection models, including data preprocessing, network architecture selection, and the training process. Additionally, it addresses key considerations such as feature extraction, model optimization, and the labeling process.
3.3.1 Choice of Neural Network Architecture
@@ 3.3.2 Feature Extraction
3.3.3 Data Analysis and Preparation
@@
Figure 6: Fall activities recorded by the left smartwatch of ID 3: $\text{FALL}_{1}$, $\text{FALL}_{2}$, $\text{FALL}_{3}$, and $\text{FALL}_{5}$, respectively.
Figure 7: Fall activities recorded by the left smartwatch of ID 3: $\text{ADL}_{1}$, $\text{ADL}_{2}$, $\text{ADL}_{3}$, $\text{ADL}_{4}$, $\text{ADL}_{5}$ and $\text{ADL}_{7}$, respectively.
There is a difference in the magnitude of activities that involve walking and running with and without a rifle ($\text{ADL}_{2}$, $\text{ADL}_{3}$, $\text{OM}_{1}$, and $\text{OM}_{2}$), with the magnitude of activities with a rifle being lower (Figure 8). This is because walking with a rifle is an engagement walk, while running is a high-speed engagement activity, not exactly a walk or a run. Among the activities that involve falls with and without a rifle, the patterns are similar, but in many cases, the magnitude of falls with a rifle is lower (Figure 8). Generally, the fall with or without the use of a rifle usually occurs in the first five seconds, followed by a stable pattern with an average acceleration magnitude of approximately 10 in the final five seconds, corresponding to the inactivity of the fallen body on the ground. Thus, the time window defined for fall detection is five seconds, corresponding, on average, to 450 observation points for the smartwatch and 1025 observation points for the smartphone. For both linear and angular acceleration, these averages are similar due to the data collection frequencies being close.
@@
Figure 8: Fall activities recorded by the left smartwatch of ID 3: $\text{ADL}_{12}$, $\text{ADL}_{12}$_withRifle, $\text{FALL}_{3}$ and $\text{FALL}_{3}$_withRifle, respectively.
Activities with data collected during trials of ten seconds or more, except for transition activities and $\text{ADL}_{5}$, $\text{ADL}_{6}$, and $\text{ADL}_{1}$5 activities related to going up and down stairs, exhibit a similar pattern in the magnitude variation over the activity duration. This allowed these activities to be segmented into more than one five-second set. For activities involving stairs, the set size was defined using only the first five seconds of the activity, discarding the rest of the collected data, as the staircase where the activities were collected was short, with only ten steps, and in most of the collected activities, after five seconds, the volunteer military personnel had already completed the ascent or descent of the stairs.
@@
In the transition activities to the shooting position ($\text{OM}_{3}$ to $\text{OM}_{8}$), such as the $\text{OM}_{4}$ activity, which is the transition from walking to the kneeling shooting position, the point of interest is the pattern caused by the state change, to verify with the neural network if, in any way, the activity is confused with a fall activity. Therefore, in these activities, only one five-second observation vector (480 and 1025) was selected and represents the moment of state change from one activity to another. In this way, the observation point where the highest magnitude peak occurs was identified, and the 240 observations before the peak and the 240 observations after the peak were selected to construct the data vector presented to the neural network, except for cases where it was not possible to distribute equally.
Fall activities and transition activities to the prone shooting position ($\text{OM}_{6}$, $\text{OM}_{7}$, and $\text{OM}_{8}$) exhibit similar patterns (Figure 9 because both involve the body hitting the ground, which can potentially confuse the neural network. However, as the first stage of the Soldier of the Future project aims to detect falls for subsequent classification if the fall occurred due to injuries or not, an adopted approach was to consider these transition activities as falls. Each experimental pipeline was executed with labels considering these activities as falls and, in another approach, considering them as non-fall activities.
Figure 9: Activities recorded by the left smartwatch of ID 3: $\text{FALL}_{3}$ and $\text{OM}_{6}$, respectively.
@@
For the execution of all experimental pipelines (Section 3.3.4), the pre-processed data for training the neural networks was divided into three sets: training, testing, and validation. The test and validation sets each correspond to 20% of the samples, while the training set corresponds to 60% of the samples. Bayesian optimization of the hyperparameters (Section 3.3.5) was performed using the training and validation data. The test set was not presented to the network during optimization or subsequent training, and was used only for making predictions and evaluating the models’ generalization performance.
@@
3.3.4 Experimental Pipelines and Labeling Approach